OpenStudy (anonymous):

Helena deposited $89 into a savings account for which interest is compounded monthly. According to the Rule of 72, what interest rate will cause her money to double in approximately 38 years? A.2.3% B.1.9% C.0.4% D.0.5%

OpenStudy (anonymous):

You have something like: 178 = 89 (1+i)^n 178 is 89 * 2 the double of money i: interest by month n: number of months n=38*12=456 now you have: 178 = 89*(1+i)^456 178/89 = (1+i)^456 2=(1+i)^456 using Log: Log(2) = 456*Log(1+i) 0.301 = 456*Log(1+i) 0.301/456 = Log(1+i) using 10^x you get: 10^(0.301/456) = 1+i solving for i you get i=0.152% i by 12 = 1.825 compounded monthly aprox. ->1.9%